# spacing=nonfrench

### v2

@@ 4.1 Experimental setup
Figure 5: Accuracy comparisons on ${\mathcal{D}_{\rm priv}}$ with calibrated $\epsilon$ values: SIDP versus MobileNetV2(${\mathcal{M}_{\rm L}}$) trained using LDPKiT-Rand and LDPKiT-Sup.
(a) Base noise with $\epsilon=1.25$
(b) Base noise with $\epsilon=1.5$
Figure 6: Comparison of LDPKiT-Rand’s inference accuracies with different $\epsilon$ values for the post-processing layer of LDP noise versus LDPKiT-Sup.
4.2 RQ1: LDPKiT’s Utility Recovery on ${\mathcal{D}_{\rm priv}}$
@@
In this section, we evaluate whether LDPKiT improves prediction accuracy on ${\mathcal{D}_{\rm priv}}$ compared to directly using ${\mathcal{M}_{\rm R}}$’s noisy labels on ${\mathcal{D}_{\rm protected}}$ (i.e., SIDP). To simulate a realistic scenario, we set $|{\mathcal{D}_{\rm priv}}|$ to 1,500, representing a practical data size a user might own. The entire ${\mathcal{D}_{\rm cand}}$ is used as ${\mathcal{D}_{\rm infer}}$ to minimize variations from random subset splits. To answer RQ1 quantitatively, we record the final accuracies of ${\mathcal{M}_{\rm L}}$ on the original ${\mathcal{D}_{\rm priv}}$ at the last epoch of training and ${\mathcal{M}_{\rm R}}$’s SIDP accuracies on ${\mathcal{D}_{\rm protected}}$. We compare the performance of SIDP and LDPKiT with different noise generation mechanisms in Figures 4 and 5. We also tabulate the numerical accuracies on ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm val}}$ in Tables 56 and 7 in Appendix D.1. The results demonstrate that the extracted ${\mathcal{M}_{\rm L}}$ with LDPKiT, regardless of the noise generation mechanism, can almost always achieve higher prediction accuracies on ${\mathcal{D}_{\rm priv}}$ compared to SIDP. For instance, when $\epsilon=2.0$, ResNet-152 (${\mathcal{M}_{\rm R}}$)’s SIDP accuracy on SVHN is only 21.07%, whereas both LDPKiT-Rand and LDPKiT-Sup can recover the inference accuracy of the trained ResNet-18 (${\mathcal{M}_{\rm L}}$) to approximately 85%. However, the improvement is not always significant and largely depends on the specific noise generation mechanism of LDPKiT. When employing LDPKiT-Rand, ResNet-18 achieves only 68.52% and 58.07% accuracy on Fashion-MNIST with $\epsilon$ values of 2.0 and 1.5, respectively, and MobileNetV2’s accuracies fall below 60%. On PathMNIST, ResNet-18’s performance is also inadequate, and MobileNetV2’s accuracy with LDPKiT-Rand is even lower than SIDP.
@@
From the two right-most bars in each column of Figures 4 and 5, we also observe that the local models learn more effectively from ${\mathcal{D}_{\rm infer}}$ created by superimposing noised data points within ${\mathcal{D}_{\rm priv}}$, i.e., LDPKiT-Sup, rather than applying arbitrary random noise as in LDPKiT-Rand. For instance, LDPKiT-Sup improves ResNet-18’s accuracy to 81.78% on Fashion-MNIST when $\epsilon$ is 2.0, compared to 28.69% with SIDP and 68.52% with LDPKiT-Rand. Similarly, ResNet-18’s accuracy is recovered to 82.11% on PathMNIST when $\epsilon$ is 10.0, far exceeding the 28.74% SIDP and 43.90% LDPKiT-Rand accuracies. Furthermore, we notice that ResNet-18 and MobileNetV2’s inference accuracies are consistently over or around 80% on SVHN when SIDP accuracy ranges between 11.87% and 21.07%. This demonstrates that LDPKiT-Sup can provide better privacy with essentially minimal loss of accuracy. Overall, the experimental results confirm that LDPKiT effectively recovers most of the utility loss caused by LDP noise under strictly privacy-preserving noise levels.
@@
Note that the base noise and post-processing noise in LDPKiT-Rand are set to the same noise level (i.e., $\epsilon$ value) by default, as described in Section 3.2. While it may seem intuitive to adjust LDPKiT-Rand by lowering the noise level of the post-processing layer (i.e., increasing $\epsilon$ for the second layer of noise) to achieve higher accuracy, empirical results suggest that maintaining the same $\epsilon$ value for the post-processing noise as the base noise is more efficient and accuracy-preserving. To explore the optimal setting of LDPKiT-Rand, we test on SVHN with ResNet-18 (${\mathcal{M}_{\rm L}}$), setting base noise $\epsilon_{1}$ to 1.5 and 1.25 with varying $\epsilon_{2}$ for post-processing noise. In Figure 6, the left-most green bar represents LDPKiT-Sup’s inference accuracy, showing consistent outperformance of LDPKiT-Rand regardless of $\epsilon$ combinations. The orange bar indicates LDPKiT-Rand’s accuracy when $\epsilon_{1}=\epsilon_{2}$. The results reveal that deviating the post-processing noise significantly from the base noise reduces LDPKiT-Rand’s utility. Although it seems like $\epsilon_{2}=2.0$ exhibits the optimal accuracy, it is worth noting that the combinations of $\epsilon_{1}=1.5,\epsilon_{2}=2.0$ and $\epsilon_{1}=1.25,\epsilon_{2}=2.0$ both achieve lower accuracies than the 84.86% obtained when $\epsilon_{1}=\epsilon_{2}=2.0$ (See Figure 4(b)). It is also empirically challenging for the user to determine the second $\epsilon$ before inference. Therefore, aligning the $\epsilon$ of the post-processing noise with the base noise is recommended for optimal efficiency and practicality.
@@
Figure 7: Latent space plots of Fashion-MNIST class triplets (C0-T-shirt/top , C2-pullover, C9-ankle boot) and privacy-protected noisy data clusters generated with LDPKiT-Rand and LDPKiT-Sup ($\epsilon=2.0$).
Table 1: Euclidean distances between centroids of clusters shown in Figure 7 on Fashion-MNIST.
Figure Strategy Class(es) $d(C_{\text{N}},C_{0})$ $d(C_{\text{N}},C_{2})$ $d(C_{\text{N}},C_{9})$
7(a) LDPKiT-Sup (0,0) 1.9835 2.5191 3.2017
LDPKiT-Rand 0 2.1195 2.4666 2.9169
7(b) LDPKiT-Sup (9,2) 3.1652 2.3538 2.0524
LDPKiT-Rand 9 3.0937 2.5595 1.9333
7(c) LDPKiT-Sup (2,2) 2.9662 1.7351 2.9171
LDPKiT-Rand 2 2.7084 1.9433 2.7115
Table 2: Frequency with which LDPKiT-Rand is more divergent from the target class triplets compared to LDPKiT-Sup.
Dataset Number of triplets with DR > 1.0 Total number of triplets Frequency
@@
DR refers to the KL Divergence Ratio introduced in Section 4.3 Definition 4.2.
@@
Table 3: Sensitivity Analysis on $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ with ResNet-18 (${\mathcal{M}_{\rm L}}$).
@@
4.3 RQ2: Latent Space Analysis
In Section 4.2, we quantitatively demonstrate that LDPKiT enhances inference accuracy on ${\mathcal{D}_{\rm priv}}$ compared to SIDP, with LDPKiT-Sup consistently outperforming the LDPKiT-Rand noise generation algorithms. We empirically demonstrate that ML models trained with LDPKiT-Sup—including ResNet-18 and MobileNetV2—achieve higher accuracies on ${\mathcal{D}_{\rm priv}}$ and ${\mathcal{D}_{\rm val}}$ than with LDPKiT-Rand. However, it is less apparent why LDPKiT-Sup is better than LDPKiT-Rand qualitatively. We hypothesize that the synthetic data points generated by LDPKiT-Sup are closer to the distribution of the original data points from target classes compared to LDPKiT-Rand, enabling more effective knowledge transfer about how ${\mathcal{M}_{\rm R}}$ classifies points in the latent space, including potentially better extraction of the decision boundaries. To evaluate our hypothesis, we plot and analyze the latent space of ${\mathcal{D}_{\rm infer}}$. To replicate our framework in latent space, we first train a variational auto-encoder (VAE) on the data samples with $\epsilon$-LDP noise from three random classes of the datasets. We pick $\epsilon$ values of 2.0 for Fashion-MNIST and SVHN, and 7.0 for PathMNIST. Other hyperparameter choices are discussed in Appendix B. Our training objective is to visualize three distinct clusters representing the original noise-free data samples corresponding to the three target classes. To achieve this, we replace the conventional reconstruction loss with the triplet margin loss. Using the trained VAE model, we further generate and visualize two additional clusters: noisy samples produced by LDPKiT-Sup based on two of the three classes, and noisy samples produced by LDPKiT-Rand based on one class. We then compute the Euclidean distances from these noisy clusters (i.e., LDPKiT-Rand and LDPKiT-Sup) to the clean class clusters. This analysis allows us to evaluate whether LDPKiT-Sup offers a learning advantage over LDPKiT-Rand, particularly through the perspective of the latent space representation.
Figure 7 presents an example of Fashion-MNIST’s latent space visualization. Clusters of Class 0, Class 2 and Class 9 represent the original (noise-free) data points from those target classes. For instance, the cluster superimposed on Class 9 and Class 2 contains noised superimposed data points generated with LDPKiT-Sup from Class 2 and Class 9. The cluster randomized on Class 0 shows noised data points generated with LDPKiT-Rand on the target Class 0. The key observation is that the data distribution of the noisy cluster created with LDPKiT-Sup is closer to the distribution of the target class(es) than that of LDPKiT-Rand. For instance, LDPKiT-Sup’s cluster is closer to the cluster of the target class (i.e., Class 0 or Class 2) than LDPKiT-Rand’s cluster in Figure 7(a) and Figure 7(c). In Figure 7(b), LDPKiT-Sup’s cluster is more spread out and has more overlapping area with the clusters of Class 2 and Class 9, whereas LDPKiT-Rand’s cluster remains more static and less representative, sitting in the middle of the clusters of Class 0, 2 and 9. Additional latent space analyses for triplet examples across other datasets are provided in Appendix D.4.
Table 1 provides the numerical Euclidean distances between the centroids of each cluster pairs from Figure 7. Here, the distance values further validate the observations from the plots. We note, however, that the Euclidean distance between centroids is not always an ideal measure of similarity between clusters. For example, as shown in Figure 7(b), the cluster generated by LDPKiT-Sup visually contains more overlapping data points with the Class 9 cluster compared to LDPKiT-Rand. However, since LDPKiT-Sup incorporates points from both Class 2 and Class 9, its distribution is more dispersed than LDPKiT-Rand which only contains data points from Class 9. Consequently, the centroid of LDPKiT-Rand is slightly closer to the centroid of the Class 9 cluster.
Definition 4.1.
KL Divergence. The KL Divergence of two probability distributions $P$ and $Q$ is defined as:
$D_{\text{KL}}(P\,||\,Q)=\sum_{x}P(x)\log\frac{P(x)}{Q(x)}$
In our case, we denote the combination of three target clusters as the target triplet cluster, $C_{T}$, consisting of all the data points in the triplet classes. The cluster generated by LDPKiT-Sup, denoted as $C_{S}$, consists of all superimposed combinations of two target classes, while LDPKiT-Rand’s generated cluster, $C_{R}$, consists of noisy data points generated on all three triplet classes. For a fair comparison, $C_{S}$ and $C_{R}$ contain the same number of points.
Definition 4.2.
where $\bar{C}_{S}$, $\bar{C}_{R}$ and $\bar{C}_{T}$ are normalized distributions that sum to 1. $\text{DR}(\bar{C}_{S},\bar{C}_{R})>1$ represents that $C_{S}$ is less divergent from $C_{T}$ than $C_{R}$ within this triplet class setting.
@@ 4.3 RQ2: Latent Space Analysis
Nuances may occur between different triplets. To evaluate the average-case scenario, we collect statistics across 130 triplets from SVHN, Fashion-MNIST, and PathMNIST and present the results in Table 2. The analysis indicates that LDPKiT-Sup generates clusters that are less divergent from the target triplet cluster, as measured by KL Divergence, in approximately 68.56% of all triplet cases, compared to LDPKiT-Rand. The finding that LDPKiT-Rand tends to diverge from the target distribution more frequently than LDPKiT-Sup match with our previous observations of higher accuracy achieved by LDPKiT-Sup and the insights gained from the latent space analysis. However, it is worth noting that the relationship between accuracy improvement and KL divergence can be complicated.
(a) PathMNIST
(b) Fashion-MNIST
(c) SVHN
Figure 9: Impact of $\epsilon$ values on ResNet-18 (${\mathcal{M}_{\rm L}}$)’s accuracies on ${\mathcal{D}_{\rm priv}}$ versus ${\mathcal{D}_{\rm val}}$ with LDPKiT-Sup.
(a) PathMNIST
(b) Fashion-MNIST
(c) SVHN
Figure 10: Impact of $\epsilon$ values on ResNet-18 (${\mathcal{M}_{\rm L}}$)’s accuracies on ${\mathcal{D}_{\rm priv}}$ versus ${\mathcal{D}_{\rm val}}$ with LDPKiT-Rand.
4.4 RQ3: Sensitivity Analysis on the Impact of $|{\mathcal{D}_{\rm infer}}|$ and $|{\mathcal{D}_{\rm priv}}|$ on LDPKiT
@@
To study the privacy leakage more thoroughly, we investigate how many private data samples ($|{\mathcal{D}_{\rm priv}}|$) are needed for ${\mathcal{D}_{\rm infer}}$ construction and how many queries ($|{\mathcal{D}_{\rm infer}}|$) are needed to train a ${\mathcal{M}_{\rm L}}$ that achieves a reasonable accuracy on ${\mathcal{D}_{\rm priv}}$. As shown in Table 3, we generate several possible ${\mathcal{D}_{\rm infer}}$ datasets from ${\mathcal{D}_{\rm priv}}$ with varying sizes to examine the effect of $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ on ${\mathcal{M}_{\rm L}}$’s accuracy. For example, when ${\mathcal{D}_{\rm priv}}$ contains 125 points, the largest ${\mathcal{D}_{\rm infer}}$ that can be created has 15,500 points (i.e., $|{\mathcal{D}_{\rm priv}}|\cdot(|{\mathcal{D}_{\rm priv}}|-1)$). If $|{\mathcal{D}_{\rm priv}}|$ is increased to 250, $|{\mathcal{D}_{\rm infer}}|$ can still be set to 15,500 by random selection or 62,250 at maximum. For variable control in analyzing LDPKiT’s sensitivity to $|{\mathcal{D}_{\rm priv}}|$, an example setting can be fixing $|{\mathcal{D}_{\rm infer}}|$ at 15,500 and comparing ${\mathcal{M}_{\rm L}}$’s accuracies when $|{\mathcal{D}_{\rm priv}}|$ is 125 versus when it is 250. Similarly, we can assess the impact of $|{\mathcal{D}_{\rm infer}}|$ by fixing $|{\mathcal{D}_{\rm priv}}|$ at 250 and comparing the ${\mathcal{M}_{\rm L}}$’s accuracy for $|{\mathcal{D}_{\rm infer}}|=15,500$ and $|{\mathcal{D}_{\rm infer}}|=62,250$. To identify trends, we exhaustively test some possible pairs of $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ for $|{\mathcal{D}_{\rm priv}}|$ ranging from 125 to 1.5k and record ResNet-18 (${\mathcal{M}_{\rm L}}$)’s accuracies in each scenario.
@@
As shown in Table 3, LDPKiT exhibits low sensitivity to variations in $|{\mathcal{D}_{\rm priv}}|$. For instance, the accuracies in the cases with fixed $|{\mathcal{D}_{\rm infer}}|$ and varying $|{\mathcal{D}_{\rm priv}}|$ are similar across the three benchmarks, e.g., on Fashion-MNIST with $\epsilon=2.0$, LDPKiT-Sup’s maximum accuracy difference among different $|{\mathcal{D}_{\rm priv}}|$ settings is below 5% when $|{\mathcal{D}_{\rm infer}}|=15,500$ and is around 0.69% when $|{\mathcal{D}_{\rm infer}}|=500k$. It also suggests that as long as $|{\mathcal{D}_{\rm infer}}|$ is reasonable, the user can generate queries with a smaller $|{\mathcal{D}_{\rm priv}}|$, which further reduces privacy risks. There is one case, however ($|{\mathcal{D}_{\rm infer}}|=250k$), where increasing $|{\mathcal{D}_{\rm priv}}|$ improves ${\mathcal{M}_{\rm L}}$’s accuracies. This is likely because a larger training set increases the probability of including more representative and diverse data points that benefit training. However, this improvement is not guaranteed.

### v3

@@ IV-A Experimental setup
Figure 5: Inference accuracy comparisons of LDPKiT-Rand on $\mathcal{D}_{\rm priv}$} with base $\epsilon=1.5$ and various post-processing $\epsilon$.
IV-B RQ1: LDPKiT’s Utility Recovery on $\mathcal{D}_{\rm priv}$}
In this section, we evaluate whether LDPKiT improves prediction accuracy on $\mathcal{D}_{\rm priv}$} compared to directly using $\mathcal{M}_{\rm R}$}’s returned labels on $\mathcal{D}_{\rm protected}$} (i.e., SIDP). To simulate a realistic scenario, we set $|{\mathcal{D}_{\rm priv}}|$ to 1,500, representing a practical data size a user might own. For variable control, we use the entire $\mathcal{D}_{\rm infer}$} with the size of $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$ for querying $\mathcal{M}_{\rm R}$} and training $\mathcal{M}_{\rm L}$} to minimize variations from random subset splits.
When no privacy protection exists and thus no noise is added, $\mathcal{M}_{\rm R}$}’s average accuracies on noise-free $\mathcal{D}_{\rm priv}$} of Fashion-MNIST, SVHN and PathMNIST are 93.33%, 94.73% and 86.2%, respectively. The accuracies on $\mathcal{D}_{\rm val}$} are 93.22%, 96.30% and 81.24%, respectively.
@@ IV-B RQ1: LDPKiT’s Utility Recovery on $\mathcal{D}_{\rm priv}$}
To answer RQ1 quantitatively, we record the last epoch’s accuracies of $\mathcal{M}_{\rm L}$} on the original $\mathcal{D}_{\rm priv}$} and $\mathcal{M}_{\rm R}$}’s SIDP accuracies on $\mathcal{D}_{\rm protected}$}. We compare SIDP and LDPKiT in Figure 4. We also tabulate the numerical accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} in Tables VVI and VII in Appendix A-D1. The results demonstrate that $\mathcal{M}_{\rm L}$} with LDPKiT can almost always achieve higher prediction accuracies on $\mathcal{D}_{\rm priv}$} compared to SIDP (the leftmost blue bar). For instance, when $\epsilon=2.0$, ResNet-152 ($\mathcal{M}_{\rm R}$})’s SIDP accuracy on SVHN is only 21.07%, whereas both LDPKiT-Rand and LDPKiT-Sup can recover the inference accuracy of the trained ResNet-18 ($\mathcal{M}_{\rm L}$}) to approximately 85%. However, the improvement is not always significant using LDPKiT-Rand. When employing LDPKiT-Rand, ResNet-18 achieves accuracies of only 68.52% and 58.07% on Fashion-MNIST with $\epsilon$ values of 2.0 and 1.5, respectively, and MobileNetV2’s accuracies fall below 60%. On PathMNIST, ResNet-18’s performance is also inadequate, and MobileNetV2’s accuracies with LDPKiT-Rand are even lower than SIDP. The results match our expectation that data generated by simply adding another layer of random noise is not representative enough for effective training. In contrast, the evaluation results of LDPKiT-Sup reported in parallel are constantly better than LDPKiT-Rand.
@@
From the LDPKiT-Rand and LDPKiT-Sup bars in Figure 4, we observe that the local models learn more effectively using $\mathcal{D}_{\rm infer}$} created by superimposing noised $\mathcal{D}_{\rm priv}$} data points, i.e., LDPKiT-Sup, rather than applying arbitrary random noise as in LDPKiT-Rand. For instance, LDPKiT-Sup improves ResNet-18’s accuracy to 81.78% on Fashion-MNIST when $\epsilon$ is 2.0, compared to 28.69% with SIDP and 68.52% with LDPKiT-Rand. Similarly, ResNet-18’s accuracy is recovered to 82.11% on PathMNIST when $\epsilon$ is 10.0, far exceeding the 28.74% SIDP and 43.90% LDPKiT-Rand accuracies. Notably, despite the low SIDP baseline accuracy (11.87%–21.07%), LDPKiT-Sup achieves comparable accuracy even under more stringent privacy settings, consistently reaching over 80% accuracy on SVHN with both ResNet-18 and MobileNetV2. This demonstrates that LDPKiT-Sup provides stronger privacy guarantees without sacrificing utility. Overall, the experimental results confirm that LDPKiT effectively recovers most of the utility loss caused by LDP noise under strictly privacy-preserving noise levels.
@@
Note that the base noise and the post-processing noise in LDPKiT-Rand are set to the same noise level (i.e., $\epsilon$ value) by default, as described in Section III-B. While it may seem intuitive to adjust LDPKiT-Rand by lowering the noise level of the post-processing layer (i.e., increasing $\epsilon$ for the second layer of noise) to achieve higher accuracy, empirical results suggest that maintaining the same $\epsilon$ value for the post-processing noise as the base noise is more efficient and accuracy-preserving. To explore the optimal configuration of LDPKiT-Rand, we conduct experiments on SVHN using a ResNet-18 model ($\mathcal{M}_{\rm L}$}), fixing the base noise $\epsilon_{1}$ at 1.5 and varying the post-processing noise $\epsilon_{2}$. In Figure 5, the left-most green bar represents LDPKiT-Sup’s inference accuracy, consistently outperforming LDPKiT-Rand across all the $\epsilon$ combinations. The orange bar (the seventh from left in Figure 5) indicates LDPKiT-Rand’s accuracy when $\epsilon_{1}=\epsilon_{2}$. The results show that deviating the post-processing noise significantly from the base noise reduces LDPKiT-Rand’s utility (e.g., inference accuracy drops to random guessing when $\epsilon>5$).
@@
Figure 6: Latent space plots of Fashion-MNIST class triplets (C0-T-shirt/top , C2-pullover, C9-ankle boot) and privacy-protected noisy data clusters generated with LDPKiT-Rand and LDPKiT-Sup ($\epsilon=2.0$). The observation is that the data clusters generated with LDPKiT-Sup are constantly closer to the target clusters than the clusters generated with LDPKiT-Rand.
TABLE I: Euclidean distances between centroids of clusters shown in Figure 6 on Fashion-MNIST, demonstrating that LDPKiT-Sup almost always generate data distribution that is closer to the target distribution than LDPKiT-Rand.
Figure Strategy Class(es) $d(C_{\textbf{\text{N}}},C_{0})$ $d(C_{\textbf{\text{N}}},C_{2})$ $d(C_{\textbf{\text{N}}},C_{9})$
6(a) Sup (0,0) 1.9835 2.5191 3.2017
Rand 0 2.1195 2.4666 2.9169
6(b) Sup (9,2) 3.1652 2.3538 2.0524
Rand 9 3.0937 2.5595 1.9333
6(c) Sup (2,2) 2.9662 1.7351 2.9171
Rand 2 2.7084 1.9433 2.7115
$d(C_{\text{N}},C_{X})$ is the Euclidean distance between the Noisy and Class X cluster centroids.
TABLE II: Frequency with which LDPKiT-Rand is more divergent from the target class triplets than LDPKiT-Sup when measuring with KL Divergence Ratio, demonstrating that LDPKiT-Sup produces more meaningful data samples than LDPKiT-Rand.
Dataset
Number of triplets
with DR > 1.0
Total number
of triplets
Frequency
@@
DR refers to the KL Divergence Ratio introduced in Section IV-C Definition IV.2.
@@
IV-C RQ2: Latent Space Analysis
In Section IV-B, we empirically demonstrate that LDPKiT enhances inference accuracy on $\mathcal{D}_{\rm priv}$} compared to SIDP, with LDPKiT-Sup consistently outperforming LDPKiT-Rand. However, it is not apparent why LDPKiT-Sup is better than LDPKiT-Rand qualitatively. We hypothesize that the synthetic data points generated by LDPKiT-Sup are closer to the distribution of the original data points from target classes compared to LDPKiT-Rand, enabling more effective knowledge transfer by generating data points with higher quality, which leads to potentially better extraction of the decision boundaries. To evaluate our hypothesis, we plot and analyze the latent space of $\mathcal{D}_{\rm infer}$}. To replicate our framework in latent space, we first train a variational auto-encoder (VAE) on the data samples with $\epsilon$-LDP noise from three random classes of the datasets. We pick $\epsilon$ values of 2.0 for Fashion-MNIST and SVHN, and 7.0 for PathMNIST. Other hyperparameter choices are discussed in Appendix A-C. Our training objective is to visualize three distinct clusters representing the original noise-free data samples corresponding to the three target classes. To achieve this, we replace the conventional reconstruction loss with the triplet margin loss. Using the trained VAE model, we further generate and visualize two additional clusters: noisy samples produced by LDPKiT-Sup based on two of the three classes, and noisy samples produced by LDPKiT-Rand based on one class. We then compute the Euclidean distances from these noisy clusters (i.e., LDPKiT-Rand and LDPKiT-Sup) to the clean class clusters. This analysis allows us to evaluate whether LDPKiT-Sup offers a learning advantage over LDPKiT-Rand, particularly from the perspective of the latent space representation.
Figure 6 presents an example of Fashion-MNIST’s latent space visualization. Clusters of Class 0, Class 2 and Class 9 represent the original (noise-free) data points from those target classes. For instance, the cluster superimposed on Class 9 and Class 2 contains noisy superimposed data points generated with LDPKiT-Sup from Classes 9 and 2. The cluster randomized on Class 0 shows noisy data points generated with LDPKiT-Rand on the target Class 0. The key observation is that the data distribution of the noisy cluster created with LDPKiT-Sup is closer to the distribution of the target class(es) than that of LDPKiT-Rand. For instance, LDPKiT-Sup’s cluster is closer to the cluster of the target class (e.g., Class 0 or Class 2) than LDPKiT-Rand’s cluster in Figure 6(a) and Figure 6(c). In Figure 6(b), LDPKiT-Sup’s cluster is more spread out and has more overlapping area with the clusters of Class 2 and Class 9, whereas LDPKiT-Rand’s cluster remains more static and less representative, sitting in the middle of the clusters of Class 0, 2 and 9. Additional latent space analyses for triplet examples across other datasets are provided in Appendix A-D4.
Table I provides the numerical Euclidean distances between the centroids of each cluster pairs from Figure 6. Here, the distance values further validate the observations from the plots. We note, however, that the Euclidean distance between centroids is not always an ideal measure of similarity between clusters. For example, as shown in Figure 6(b), the cluster generated by LDPKiT-Sup visually contains more overlapping data points with the Class 9 cluster compared to LDPKiT-Rand. However, since LDPKiT-Sup incorporates points from both Class 2 and Class 9, its distribution is more dispersed than LDPKiT-Rand, which only contains data points from Class 9. Consequently, the centroid of LDPKiT-Rand is slightly closer to the centroid of the Class 9 cluster.
Definition IV.1.
Kullback-Leibler (KL) Divergence. The KL Divergence of two probability distributions $P$ and $Q$ is defined as:
$D_{\text{KL}}(P\,||\,Q)=\sum_{x}P(x)\log\frac{P(x)}{Q(x)}$
In our case, we denote the combination of three target clusters as the target triplet cluster, $C_{T}$, consisting of all the data points in the triplet classes. The cluster generated by LDPKiT-Sup, denoted as $C_{S}$, consists of all the superimposed combinations of two target classes, while LDPKiT-Rand’s generated cluster, $C_{R}$, consists of noisy data points generated on all three triplet classes. For a fair comparison, $C_{S}$ and $C_{R}$ contain the same number of points.
Definition IV.2.
where $\bar{C}_{S}$, $\bar{C}_{R}$ and $\bar{C}_{T}$ are the normalized distributions of $C_{S}$, $C_{R}$ and $C_{T}$. $\text{DR}(\bar{C}_{R},\bar{C}_{S})>1.0$ means that $\bar{C}_{S}$ is less divergent from $\bar{C}_{T}$ than $\bar{C}_{R}$ in this triplet class setting (i.e., $C_{S}$ is a better representation of $C_{T}$ than $C_{R}$).
TABLE III: Sensitivity analysis of LDPKiT on different dataset sizes, $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$, with ResNet-18 ($\mathcal{M}_{\rm L}$}).
@@ IV-C RQ2: Latent Space Analysis
Figure 7: Comparison of ResNet18 ($\mathcal{M}_{\rm L}$})’s accuracies on $\mathcal{D}_{\rm priv}$} and $\mathcal{D}_{\rm val}$} with $|{\mathcal{D}_{\rm priv}}|=1.5k$ and various $|{\mathcal{D}_{\rm infer}}|$. The results show that LDPKiT-Sup demonstrates higher accuracy when $|{\mathcal{D}_{\rm infer}}|$ is larger.
@@
Nuances may occur between different triplets. To evaluate the average-case scenario, we collect statistics across 130 triplets from SVHN, Fashion-MNIST, and PathMNIST and present the results in Table II. The analysis indicates that in approximately 68.56% of all the triplet cases, LDPKiT-Sup generates clusters that are less divergent from the target triplet cluster compared to LDPKiT-Rand, measured by KL Divergence. The finding that LDPKiT-Rand tends to diverge from the target distribution more frequently than LDPKiT-Sup matches our previous observations of higher accuracy achieved by LDPKiT-Sup and the insights gained from the latent space analysis. However, it is worth noting that the relationship between accuracy improvement and KL divergence can be complicated and requires further statistical analysis.
IV-D RQ3: Sensitivity Analysis of the Impact of $|{\mathcal{D}_{\rm infer}}|$ and $|{\mathcal{D}_{\rm priv}}|$ on LDPKiT
@@
To study the trade-off among privacy ($|{\mathcal{D}_{\rm priv}}|$), cost ($|{\mathcal{D}_{\rm infer}}|$) and accuracy more thoroughly, we investigate how many private data samples ($|{\mathcal{D}_{\rm priv}}|$) are needed for augmentation and how many queries ($|{\mathcal{D}_{\rm infer}}|$) are needed to train an $\mathcal{M}_{\rm L}$} that achieves a reasonable accuracy on $\mathcal{D}_{\rm priv}$}. As shown in Table III, we generate several possible $\mathcal{D}_{\rm infer}$} datasets from $\mathcal{D}_{\rm priv}$} with varying sizes to examine the effect of $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ on $\mathcal{M}_{\rm L}$}’s accuracy. For example, when $\mathcal{D}_{\rm priv}$} contains 125 points, the largest possible $\mathcal{D}_{\rm infer}$} can have 15,500 points (i.e., $|{\mathcal{D}_{\rm protected}}|\cdot(|{\mathcal{D}_{\rm protected}}|-1)$). If $|{\mathcal{D}_{\rm priv}}|$ is increased to 250, $|{\mathcal{D}_{\rm infer}}|$ can still be set to 15,500 by random selection or 62,250 at maximum. For variable control in analyzing LDPKiT’s sensitivity to $|{\mathcal{D}_{\rm priv}}|$, an example setting can be fixing $|{\mathcal{D}_{\rm infer}}|$ at 15,500 and comparing $\mathcal{M}_{\rm L}$}’s accuracies when $|{\mathcal{D}_{\rm priv}}|$ is 125 versus when it is 250. Similarly, we can assess the impact of $|{\mathcal{D}_{\rm infer}}|$ by fixing $|{\mathcal{D}_{\rm priv}}|$ at 250 and comparing $\mathcal{M}_{\rm L}$}’s accuracy for $|{\mathcal{D}_{\rm infer}}|=15,500$ and $|{\mathcal{D}_{\rm infer}}|=62,250$. To identify trends, we exhaustively evaluate various pairs of $|{\mathcal{D}_{\rm priv}}|$ and $|{\mathcal{D}_{\rm infer}}|$ for $|{\mathcal{D}_{\rm priv}}|$ ranging from 125 to 1.5k and record ResNet-18 ($\mathcal{M}_{\rm L}$})’s accuracies in each scenario.
@@
As shown in Table III, LDPKiT exhibits lower sensitivity to variations in $|{\mathcal{D}_{\rm priv}}|$, which means that the user can get similarly accurate predictions without using more private data (i.e., LDPKiT provides similar utility with less privacy risk). For instance, the accuracies in the cases with fixed $|{\mathcal{D}_{\rm infer}}|$ and varying $|{\mathcal{D}_{\rm priv}}|$ are similar across the three benchmarks, e.g., on Fashion-MNIST with $\epsilon=2.0$, LDPKiT-Sup’s maximum accuracy difference among different $|{\mathcal{D}_{\rm priv}}|$ settings is below 5% when $|{\mathcal{D}_{\rm infer}}|=15,500$ and is around 0.69% when $|{\mathcal{D}_{\rm infer}}|=500k$. It also suggests that as long as $|{\mathcal{D}_{\rm infer}}|$ is reasonable, the user can generate queries with a smaller $|{\mathcal{D}_{\rm priv}}|$, which further reduces privacy risks. There is one case, however ($|{\mathcal{D}_{\rm infer}}|=250k$), where increasing $|{\mathcal{D}_{\rm priv}}|$ improves $\mathcal{M}_{\rm L}$}’s accuracies. This is likely because a larger training set increases the probability of including more representative and diverse data points that benefit training. However, this improvement is not guaranteed.